# 2INFN, Laboratori Nazionali del Gran Sasso, I-67100 Assergi, Italy 3Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert Ueberle Str. 2, 69120, Heidelberg, Germany 4Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Heidelberg, Germany

### v1

@@ 2.5 Binary initial conditions
The simulation sets we analyze here are the same as those presented by Costa et al. (2023). In this work, we analyze BNSs and BHNSs for the first time, whereas Costa et al. (2023) focused on binary BHs.
We assume the following fiducial models for the IMFs and initial orbital properties of the simulated binary stars. For both Pop. II and Pop. III binary stars, we randomly sample the initial mass ratios, orbital periods and eccentricities from the distributions reported by Sana et al. (2012) and obtained from observations of Galactic O-type binary stars. For Pop. III binary stars, we randomly sample the ZAMS masses of Pop. III primaries from a log-flat IMF, as motivated by previous studies (Stacy & Bromm 2013; Susa et al. 2014; Hirano et al. 2015; Wollenberg et al. 2020; Chon et al. 2021; Tanikawa et al. 2021a; Jaura et al. 2022; Prole et al. 2022). In contrast, we randomly sample the ZAMS masses of Pop. II primaries from a Kroupa IMF (Kroupa 2001). Hereafter, we refer to our fiducial model for Pop. III and Pop. II binary stars as log1 and kro1, respectively, for consistency with previous works (Costa et al. 2023; Santoliquido et al. 2023; Mestichelli et al. 2024).
In Sections 3.1.1, 3.2.1, and 4.1, we compare our fiducial models with additional simulations that assume a variety of initial conditions, encompassing the uncertainties on both IMF and orbital properties of Pop. III binary stars. We describe the additional models in Appendix A and summarize them in Table 1. We wish to point out that in this work we do not explore the possibility of a minimum ZAMS mass smaller than $2\,{\rm M}_{\odot}$, and, as a consequence, we do not analyze the impact of stars with $m_{\rm ZAMS}<2\,{\rm M}_{\odot}$.
@@
kro1 Kroupa (2001) Flat in log S12 S12 S12
kro5 Kroupa (2001) Flat in log SB13 SB13 Thermal
lar1 Larson (1998) - S12 S12 S12
lar5 Larson (1998) - SB13 SB13 Thermal
log1 Flat in log - S12 S12 S12
log2 Flat in log - S12 SB13 Thermal
log3 - - Sorted S12 S12
log4 - - SB13 S12 Thermal
log5 Flat in log - SB13 SB13 Thermal
top1 Top heavy - S12 S12 S12
top5 Top heavy Flat in log SB13 SB13 Thermal
Table 1: Initial conditions for the stellar population models analyzed in this work. Column 1: name of the model. Column 2: IMF for the primary star. Column 3: ZAMS mass of the overall stellar population. Columns 4, 5 and 6: initial distributions for the mass ratio $q$, period $P$ and eccentricity $e$. The acronyms S12 and SB13 refer to Sana et al. (2012) and Stacy & Bromm (2013), respectively.
@@ 2.6 Formation channels
To interpret the properties of the simulated BNS and BHNS mergers, we will refer to the six formation channels introduced by Iorio et al. (2023) and defined as follows. Channel 1 involves stable mass transfer before the formation of the first compact object, followed by at least one common-envelope phase. In channel 2, the binary evolves only through stable mass transfer episodes. In both channels 3 and 4, the binary undergoes at least one common-envelope phase before the formation of the first compact object. In particular, in channel 3, the companion star retains an H-rich envelope at the time of formation of the first compact object. In channel 4, instead, the companion has been stripped of its envelope when the first compact remnant forms. The two remaining channels refer to systems that do not interact at all (channel 0), or interact only after the formation of the first compact object (channel 5). A summary of these formation channels can be found in Table 2.
@@ 5 Interaction only after formation of first remnant
Table 2: Formation channels of merging compact binaries. Column 1: number of the channel. Column 2: brief description of the formation channel. Here SMT stands for stable mass transfer, and CE stands for common envelope.
@@ 2.7 Merger rate density
We estimate the merger rate density (MRD) of compact binaries using the semi-analytic code $\mathcal{R}$ate (Santoliquido et al. 2020, 2021, 2023), which combines catalogs of simulated compact binary mergers with a model for the metal-dependent cosmic star formation rate density. Specifically, the MRD in the comoving frame is computed as:
@@
where $\psi(z^{\prime},Z)$ is the metallicity-dependent cosmic star formation rate density at redshift $z^{\prime}$ and metallicity $Z$, and ${{\rm{d}}t(z^{\prime})}/{{\rm{d}}z^{\prime}}=H_{0}^{-1}\,{}(1+z^{\prime})^{-1}\,{}[(1+z^{\prime})^{3}\Omega_{M}+\Omega_{\Lambda}{}]^{-1/2}$, where $H_{0}$ is the Hubble parameter and $\Omega_{M}$ and $\Omega_{\Lambda}$ are the matter and energy density, respectively. We adopt here the values of the cosmological parameters from Planck Collaboration et al. (2020).
The term $\mathcal{F}(z^{\prime},z,Z)$ in Equation 2 is given by:
@@
where ${{\rm{d}}\mathcal{N}(z^{\prime},z,Z)/{\rm{d}}}t(z)$ is the rate of BNS or BHNS mergers from stars with metallicity $Z$ that form at redshift $z^{\prime}$ and merge at redshift $z$, and $M_{\ast}(z^{\prime},Z)$ is the initial total stellar mass that forms at redshift $z^{\prime}$ with metallicity $Z$. We obtain ${{\rm{d}}\mathcal{N}(z^{\prime},z,Z)/{\rm{d}}}t(z)$ directly from our simulations.
Owing to the lack of direct observations of Pop. III stars, numerous models have been proposed for their $\psi(z^{\prime},Z)$. We adopt the star formation history derived from the semi-analytic model a-sloth, developed by Hartwig et al. (2022) and Magg et al. (2022), which tracks individual Pop. III and Pop. II stars and is calibrated against a range of observables from both the local and high-redshift Universe. For comparison, we also estimate the MRD of BNSs and BHNSs formed from early Pop. II stars. In this case, we adopt the star formation rate density of Pop. II stars ($Z\sim 10^{-5}-10^{-2}\,\rm Z_{\odot}$) by Liu et al., in prep., obtained adopting a revised version of the semi-analytic code a-sloth (Hartwig et al. 2024). This version is calibrated to reproduce the metallicity – stellar mass – star formation rate relation of high-redshift galaxies in recent observations of the James Webb Space Telescope (Sarkar et al. 2025).
The described star formation rate densities $\psi(z^{\prime},Z)$ are reported in Fig. 2. The adopted models for $\psi(z^{\prime},Z)$ predicted by a-sloth are valid for $z>4.5$. Therefore, it is assumed that Pop. III and Pop. II star formation at later epochs is negligible. This is supported by the finding that their star formation rate density peaks at $z\sim 10-15$, and has dropped by a factor of 5 - 50 by $z\sim 4.5$, because of metal enrichment and re-ionization. We also show the uncertainty range associated with Pop. II star formation, reflecting variations in SN-driven outflows and assumptions on the IMF (Liu et al., in prep.)
Figure 2: Star formation rate density of Pop. III (magenta dash-dotted line, Hartwig et al. 2022) and Pop. II (purple dashed line, Liu et al., in prep.) stars. The shaded purple region denotes the uncertainty range associated with Pop. II star formation history. As a comparison, the black solid line shows the star formation rate density of Pop. II-I stars by Madau & Fragos (2017).
@@ 3 Results
3.1 Black hole – neutron star binaries (BHNSs)
In this Section, we describe the properties of BHNS mergers from Pop. III and Pop. II binary stars. We first discuss the impact of the simulated Pop. III models on the MRD of BHNSs and make a comparison with the MRD of Pop. II BHNSs. We then explore the distributions of delay times, primary masses and mass ratios of BHNSs for the fiducial models of Pop. III and Pop. II star progenitors.
@@ 3.1.1 Merger rate density
Figure 3 shows the MRD of BHNSs. Here, we consider all the models summarized in Table 1 for Pop. III stars. For Pop. II stars, instead, we take into account only the fiducial model (kro1). As a comparison, we also show the total MRD of Pop. II-I BHNSs reported by Iorio et al. (2023) (see Appendix B for details).
@@
The MRD of BHNSs from Pop. III stars peaks at $z\sim 13$, where the assumed models span about three orders of magnitude $\left[\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 10^{-2}-2\,\rm Gpc^{-3}\,yr^{-1}\right]$. The position of the peak is the result of a convolution between the chosen Pop. III star formation rate density (Hartwig et al. 2022) and the predominantly short delay times of BHNS mergers (see Sec. 3.1.2 and Fig. 4). At the peak, the largest MRD is associated with model lar1 (see Table 1 and Appendix A). Model lar1 produces a $\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 1.7\,\rm Gpc^{-3}\,yr^{-1}$, approximately four times larger than the MRD peak of our fiducial model ($\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 0.4\,\rm Gpc^{-3}\,yr^{-1}$). In fact, although top-heavy IMFs tend to produce more BH and neutron star progenitors, the expansion of these massive progenitors during their evolution very often leads to premature mergers of the stellar components, inhibiting the formation of a compact binary (see Fig. 1). At $z=0$, all the Pop. III star models produce MRDs between $\sim 10^{-3}$ and $\sim 10^{-1}\,\rm Gpc^{-3}\,yr^{-1}$. In agreement with Santoliquido et al. (2023), we find that models with initial orbital periods from Sana et al. (2012) produce larger MRDs than models relying on Stacy & Bromm (2013) (log2, log5, kro5, lar5, top5). In fact, the former have smaller orbital periods than the latter, favoring BHNS mergers.
@@
The MRD of Pop. II BHNSs peaks at $z\sim 9$ reaching $\mathcal{R}_{\rm BHNS}(z\sim 9)\sim 14\,\rm Gpc^{-3}\,yr^{-1}$. At $z=0$, Pop. II BHNSs yield a MRD approximately seven times higher than that of Pop. III stars under model lar1. The MRD from Pop. II BHNSs continues to dominate over that of Pop. III stars up to $z\sim 17$, even under the most favorable assumptions for Pop. III.
@@
Finally, Pop. II-I BHNSs yield a MRD peaking at $z\sim 2.5$ ($\mathcal{R}_{\rm BHNS}(z\sim 2.5)\sim 60\,\rm Gpc^{-3}\,yr^{-1}$) due to the combined effect of star formation rate density (Madau & Fragos 2017) and dependence on metallicity (Iorio et al. 2023). Our comparison sample from Iorio et al. (2023) including Pop. I stars dominates over the contribution of Pop. II stars up to $z\sim 7$.
@@
Here, we note the seeming contradiction that the BHNS MRD from Pop. II-I stars is lower than the BHNS MRD from Pop. II stars alone at $z\sim{7-15}$. However, we remind that the compact object MRD from Pop. II-I stars reported by Iorio et al. (2023) was obtained with several different assumptions compared to the new MRDs presented here. The main differences are the maximum ZAMS mass (assumed to be only 150 M by Iorio et al. 2023), and the different cosmic star formation rate density $\psi{}(z,Z)$. Specifically, Iorio et al. (2023) obtain $\psi{}(z,Z)$ from Madau & Fragos (2017) assuming a log-normal metallicity distribution with a narrow dispersion $\sigma_{\mathrm{Z}}=0.1$ (see Fig. 2 and Appendix B), whereas here we assume self-consistent values of $\psi{}(z,Z)$ for both Pop. II and III stars directly derived from a-sloth (Hartwig et al. 2022, 2024; Liu at al., in prep.). Assuming a narrow metallicity dispersion suppresses pockets of metal-poor star formation, as already discussed by Sgalletta et al. (2024).
Table 3 reports a summary of the main results of this Section.
@@
Table 3: Merger rate density of BHNSs from Pop. III and Pop. II stars. Column 1: name of the model producing the largest merger rate density. In the case of Pop. II stars, only the fiducial model kro1 was considered. Column 2: stellar population. Column 3: MRD at $z=0$. Column 4: MRD at the peak of the Pop. III distribution. Column 5: MRD at the peak of the Pop. II distribution.
Figure 3: Merger rate density of BHNSs as a function of the redshift $z$. The colored lines show the distributions for all the simulated Pop. III models. The thick solid blue line represents our fiducial model for Pop. III stars (log1). The thick solid gray line is the fiducial model for Pop. II stars (kro1); the gray shaded area encompasses the uncertainties on galactic outflows and IMF parameters (see Sec. 2.7). Finally, the thick solid black line represents the merger rate density of Pop. II-I BHNSs estimated by Iorio et al. (2023).
@@
Table 4: Formation channels of BHNS mergers. Column 1: name of the model. Column 2: stellar population. Column 3: merger efficiency of BHNSs $\eta_{\rm BHNS}=N_{\rm BHNS,m}/M_{*}$, with $N_{\rm BHNS,m}$ the number of BHNS mergers. Column 4-9: percentage of BHNS mergers happening through channel 0-5.

### v2

@@ 2.5 Binary initial conditions
The simulation sets we analyze here are the same as those presented by costa2023. In this work, we analyze BNSs and BHNSs for the first time, whereas costa2023 focus on binary BHs.
We assumed the following fiducial models for the IMFs and initial orbital properties of the simulated binary stars. For both Pop. II and Pop. III binary stars, we randomly sampled the initial mass ratios, orbital periods, and eccentricities from the distributions reported by sana2012 and obtained from observations of Galactic O-type binary stars. For Pop. III binary stars, we randomly sampled the ZAMS masses of Pop. III primaries from a log-flat IMF, which we were motivated to do by previous studies (sb2013; susa2014; hirano2015; wollenberg2020; chon2021; tanikawa2021b; jaura2022; prole2022). In contrast, we randomly sampled the ZAMS masses of Pop. II primaries from a Kroupa IMF (kroupa2001). Hereafter, we refer to our fiducial model for Pop. III and Pop. II binary stars as log1 and kro1, respectively, for consistency with previous works (costa2023; santoliquido2023; mestichelli2024).
In Sections 3.1.1, 3.2.1, and 4.1, we compare our fiducial models with additional simulations that assume a variety of initial conditions, encompassing the uncertainties on both IMF and orbital properties of Pop. III binary stars. We describe the additional models in Appendix A and summarize them in Table 1. Here, we do not consider stars with $m_{\rm ZAMS}<2\,{\rm M}_{\odot}$.
Table 1: Initial conditions for the stellar models [^1] Column 1: Name of the model. Column 2: IMF of the primary star. Column 3: ZAMS mass of the overall stellar population. Columns 4, 5, and 6: initial distributions for the mass ratio, $q$, period, $P$, and eccentricity, $e$. The acronyms sana2012 and sb2013 refer to sana2012 and sb2013, respectively.
@@
kro1 kroupa2001 Flat in log sana2012 sana2012 sana2012
kro5 kroupa2001 Flat in log sb2013 sb2013 Thermal
lar1 larson1998 - sana2012 sana2012 sana2012
lar5 larson1998 - sb2013 sb2013 Thermal
log1 Flat in log - sana2012 sana2012 sana2012
log2 Flat in log - sana2012 sb2013 Thermal
log3 - - Sorted sana2012 sana2012
log4 - - sb2013 sana2012 Thermal
log5 Flat in log - sb2013 sb2013 Thermal
top1 Top heavy - sana2012 sana2012 sana2012
top5 Top heavy Flat in log sb2013 sb2013 Thermal
@@ 2.6 Formation channels
To interpret the properties of the simulated BNS and BHNS mergers, we refer to the six formation channels introduced by sevn2023 and defined as follows. Channel 1 involves stable mass transfer before the formation of the first compact object, followed by at least one common-envelope phase. In channel 2, the binary evolves only through stable mass transfer episodes. In both channels 3 and 4, the binary undergoes at least one common-envelope phase before the formation of the first compact object. In particular, in channel 3, the companion star retains an H-rich envelope at the time of formation of the first compact object. In channel 4, instead, the companion has been stripped of its envelope when the first compact remnant forms. The two remaining channels refer to systems that do not interact at all (channel 0), or interact only after the formation of the first compact object (channel 5). A summary of these formation channels can be found in Table 2.
Table 2: Formation channels of merging compact binaries [^2] Column 1: Number of the channel. Column 2: Brief description of the formation channel. Here SMT stands for stable mass transfer, and CE stands for common envelope.
@@ 2.7 Merger rate density
[…]
@@
where $\psi(z^{\prime},Z)$ is the metallicity-dependent cosmic star formation rate density at redshift $z^{\prime}$ and metallicity $Z$, and ${{\rm{d}}t(z^{\prime})}/{{\rm{d}}z^{\prime}}=H_{0}^{-1}\,{}(1+z^{\prime})^{-1}\,{}[(1+z^{\prime})^{3}\Omega_{M}+\Omega_{\Lambda}{}]^{-1/2}$, where $H_{0}$ is the Hubble parameter and $\Omega_{M}$ and $\Omega_{\Lambda}$ are the matter and energy density, respectively. We adopted here the values of the cosmological parameters from aghanim2020.
The term $\mathcal{F}(z^{\prime},z,Z)$ in Equation 2 is given by
@@
where ${{\rm{d}}\mathcal{N}(z^{\prime},z,Z)/{\rm{d}}}t(z)$ is the rate of BNS or BHNS mergers from stars with metallicity $Z$ that form at redshift $z^{\prime}$ and merge at redshift $z$, and $M_{\ast}(z^{\prime},Z)$ is the initial total stellar mass that forms at redshift $z^{\prime}$ with metallicity $Z$. We obtained ${{\rm{d}}\mathcal{N}(z^{\prime},z,Z)/{\rm{d}}}t(z)$ directly from our simulations.
Owing to the lack of direct observations of Pop. III stars, numerous models have been proposed for their $\psi(z^{\prime},Z)$ (e.g., jaacks2019; skinner2020; liu2020). We adopted the star formation history derived from the semi-analytic code a-sloth, developed by hartwig2022 and magg2022, which tracks individual Pop. III and Pop. II stars and is calibrated against a range of observables from both the local and high-redshift Universe. For comparison, we also estimated the MRD of BNSs and BHNSs formed from early Pop. II stars. In this case, we adopted the star formation rate density of Pop. II stars ($Z\sim 10^{-5}-10^{-2}\,\rm Z_{\odot}$) by liu2025, obtained adopting a revised version of the semi-analytic code a-sloth (hartwig2024). This version is calibrated to reproduce the metallicity–stellar mass–star formation rate relation of high-redshift galaxies in recent observations of the James Webb Space Telescope (sarkar2025).
The described star formation rate densities, $\psi(z^{\prime},Z)$, are reported in Fig. 2. The adopted models for $\psi(z^{\prime},Z)$ predicted by a-sloth are valid for $z>4.5$. Therefore, it is assumed that Pop. III and early Pop. II star formation at later epochs is negligible. This is supported by the finding that their star formation rate density peaks at $z\sim 10-15$, and has dropped by a factor of 5 - 50 by $z\sim 4.5$, because of metal enrichment and reionization. We also show the uncertainty range associated with Pop. II star formation, reflecting variations in SN-driven outflows and assumptions about the IMF (liu2025).
Figure 2: Star formation rate density of Pop. III (dash-dotted magenta line, hartwig2022) and Pop. II (dashed purple line, liu2025) stars. The shaded purple region denotes the uncertainty range associated with Pop. II star formation history. As a comparison, the solid black line shows the star formation rate density of Pop. II-I stars by madau2017.
@@ 3 Results
3.1 Black hole—neutron star binaries (BHNSs)
In this Section, we describe the properties of BHNS mergers from Pop. III and Pop. II binary stars. We first discuss the impact of the simulated Pop. III models on the MRD of BHNSs and make a comparison with the MRD of Pop. II BHNSs. We then explore the distributions of delay times, primary masses, and mass ratios of BHNSs for the fiducial models of Pop. III and Pop. II star progenitors.
@@ 3.1.1 Merger rate density
Figure 3 shows the MRD of BHNSs. Here, we consider all the models summarized in Table 1 for Pop. III stars. For Pop. II stars, instead, we take into account only the fiducial model (kro1). As a comparison, we also show the total MRD of Pop. II-I BHNSs reported by sevn2023 (see Appendix B for details).
@@
The MRD of BHNSs from Pop. III stars peaks at $z\sim 13$, where the assumed models span about three orders of magnitude, $\left[\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 10^{-2}-2\,\rm Gpc^{-3}\,yr^{-1}\right]$. The position of the peak is the result of a convolution between the chosen Pop. III star formation rate density (hartwig2022) and the predominantly short delay times of BHNS mergers (see Sec. 3.1.2 and Fig. 4). At the peak, the largest MRD is associated with model lar1 (see Table 1 and Appendix A). Model lar1 produces $\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 1.7\,\rm Gpc^{-3}\,yr^{-1}$, which is approximately four times larger than the MRD peak of our fiducial model ($\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 0.4\,\rm Gpc^{-3}\,yr^{-1}$). In fact, although top-heavy IMFs tend to produce more BH and neutron star progenitors, the expansion of these massive progenitors during their evolution very often leads to premature mergers of the stellar components, inhibiting the formation of a compact binary (see Fig. 1). At $z=0$, all the Pop. III star models produce MRDs between $\sim 10^{-3}$ and $\sim 10^{-1}\,\rm Gpc^{-3}\,yr^{-1}$. In agreement with santoliquido2023, we find that models with initial orbital periods from sana2012 produce larger MRDs than models relying on sb2013 (log2, log5, kro5, lar5, top5). In fact, the former have smaller orbital periods than the latter, favoring BHNS mergers. In Appendix C, we repeat this analysis using different models of Pop. III star formation rate densities.
@@
The MRD of Pop. II BHNSs peaks at $z\sim 9$, reaching $\mathcal{R}_{\rm BHNS}(z\sim 9)\sim 14\,\rm Gpc^{-3}\,yr^{-1}$. At $z=0$, Pop. II BHNSs yield a MRD approximately seven times higher than that of Pop. III stars under model lar1. The MRD from Pop. II BHNSs continues to dominate over that of Pop. III stars up to $z\sim 17$, even under the most favorable assumptions for Pop. III stars.
@@
Finally, Pop. II-I BHNSs yield a MRD peaking at $z\sim 2.5$ ($\mathcal{R}_{\rm BHNS}(z\sim 2.5)\sim 60\,\rm Gpc^{-3}\,yr^{-1}$) due to the combined effect of the star formation rate density (madau2017) and the dependence on metallicity (sevn2023). Our comparison sample from sevn2023 including Pop. I stars dominates over the contribution of Pop. II stars up to $z\sim 7$.
@@
Here, we note the seeming contradiction that the BHNS MRD from Pop. II-I stars is lower than the BHNS MRD from Pop. II stars alone at $z\sim{7-15}$. However, we note that the compact object MRD from Pop. II-I stars reported by sevn2023 was obtained with several different assumptions compared to the new MRDs presented here. The main differences are the maximum ZAMS mass (assumed to be only 150 M by sevn2023), and the different cosmic star formation rate density, $\psi{}(z,Z)$. Specifically, sevn2023 obtain $\psi{}(z,Z)$ from madau2017 assuming a log-normal metallicity distribution with a narrow dispersion $\sigma_{\mathrm{Z}}=0.1$ (see Fig. 2 and Appendix B), whereas here we assume self-consistent values of $\psi{}(z,Z)$ for both Pop. II and III stars directly derived from a-sloth (hartwig2022; hartwig2024; Liu at al., in prep.). Assuming a narrow metallicity dispersion suppresses pockets of metal-poor star formation, as has already been discussed by sgalletta2025. Table 3 reports a summary of the main results of this section.
Table 3: Merger rate density of BHNSs from Pop. III and Pop. II stars [^3] Column 1: Name of the model producing the largest MRD. In the case of Pop. II stars, only the fiducial model kro1 was considered. Column 2: Stellar population. Column 3: MRD at $z=0$. Column 4: MRD at the peak of the Pop. III distribution. Column 5: MRD at the peak of the Pop. II distribution.
@@
Figure 3: Merger rate density of BHNSs as a function of the redshift, $z$. The colored lines show the distributions for all the simulated Pop. III models. The thick solid blue line represents our fiducial model for Pop. III stars (log1). The thick solid gray line is the fiducial model for Pop. II stars (kro1); the shaded gray area encompasses the uncertainties on galactic outflows and IMF parameters (see Sec. 2.7). Finally, the thick solid black line represents the MRD of Pop. II-I BHNSs estimated by sevn2023.
Table 4: Formation channels of BHNS mergers [^4] Column 1: Name of the model. Column 2: Stellar population. Column 3: Merger efficiency of BHNSs, $\eta_{\rm BHNS}=N_{\rm BHNS,m}/M_{*}$, with $N_{\rm BHNS,m}$ the number of BHNS mergers. Column 4-9: Percentage of BHNS mergers happening through channel 0-5.